Given that a positive charge Q1=+3c and a negative charge Q2=-2c,are separated by a distance of 8cm.Take permittivity constant k=9×10E9Nm2/C2.
i.find the electric field at a point p,1cm from Q2(m)
ii.what is the electric potential at a point p,5cm from Q1.
1
Expert's answer
2021-05-31T08:40:05-0400

i) The electric field at that point is the superposition of the individual charge fields. The field from "Q_1" is given as follows:


"E_1 = \dfrac{kQ_1}{r_1^2}"

where "r_1 = 8cm - 1cm = 7cm = 0.07m" is the distance from the first charge to the point.

Similarly:


"E_2 = \dfrac{kQ_2}{r_2^2}"

where "r_2 = 1cm = 0.01m" is the distance from "Q_2" to the point.

The total field is:


"E = E_1 + E_2 = \dfrac{kQ_1}{r_1^2} + \dfrac{kQ_2}{r_2^2} = k\left( \dfrac{Q_1}{r_1^2} + \dfrac{Q_2}{r_2^2} \right)\\
E = 9\times 10^9\cdot \left( \dfrac{3}{0.07^2} - \dfrac{2}{0.01^2} \right) \approx -1.7\times 10^{14}N/C"

ii) Similarly:


"E = k\left( \dfrac{Q_1}{r_1^2} + \dfrac{Q_2}{r_2^2} \right)"

where now "r_1 = 5cm = 0.05m" is the distance from "Q_1" to the point, and "r_2 = 8cm - 5cm = 0.03m" is the distance from "Q_2" to the point. Substituting the numbers, obtain:


"E = 9\times 10^9\cdot \left( \dfrac{3}{0.05^2} - \dfrac{2}{0.03^2} \right) = -9.2\times 10^{12}N/C"

Answer. i) "-1.7\times 10^{14}N/C", ii) "-9.2\times 10^{12}N/C".